on a 30 question test how many can you miss in order to receive a 70%?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

on a 30 question test how many can you miss in order to receive a 70%?
--
0.70*30 = 21
---
30-21 = 9
---
You can miss 9
==================
Cheers,
Stan H.

Answer by mananth(16949)

questions 30
required 70%
..
can miss 30 %
30 *30/100= 9
..
you can miss 9 questions
...
m.ananth@hotmail.ca